OpenStudy (anonymous):

simplify Radical 5x+21=radical 6 possible answers x=-3 x=3 x=11 none of the above

OpenStudy (anonymous):

1. Remove Radical by taking square of both sides 2. Solve for x

OpenStudy (anonymous):

Please use parenthesis... It's either one of these depending on what the radical is: \[\sqrt{5 x+21} = \sqrt{6}\] \[5 x+21 = 6 \] \[5x = -15\] \[x=-3\] OR \[\sqrt{5x}+21 = \sqrt{6}\] \[\sqrt{5} * \sqrt{x}+21 = \sqrt{6}\] \[ \sqrt{x} = \sqrt{\frac{6}{5}}-\frac{21}{\sqrt{5}} \] \[x = \frac{447-42 \sqrt(6)}{5} \]

OpenStudy (anonymous):

You chose based on what you meant, both of those are correct depending on what is having a squareroot of (5x) or (5x+21), and it makes a BIG difference as you can see.